OpenStudy (anonymous):

√11x+3(all rooted)-2x=0

OpenStudy (anonymous):

\[\sqrt{11x+3}=2x\] \[11x+3=4x^2\] \[4x^2-11x-3=0\] \[(4x+1)(x-3)=0\] \[x=3\] \[x=-\frac{1}{4}\]

OpenStudy (anonymous):

gotta check though!

OpenStudy (anonymous):

\[\sqrt{11\times 3+3}=2\times 3\] \[\sqrt{36}=6\] this one checks

OpenStudy (anonymous):

next one doesnt work so only 6